OpenStudy (anonymous):

please i need help will fan and medal for the correct answer... i need help asap

OpenStudy (anonymous):

http://prntscr.com/a5cpn7

OpenStudy (samwisegamge):

geese I know how to do it buts hard to explain

OpenStudy (anonymous):

i really just need to get done nothing bad a bout you i just

OpenStudy (anonymous):

is this the same thng as the last one???

OpenStudy (zombrexi):

y=2x+6;3x−y=6 Step 1: Solve y=2x+6for y: Step 2: Substitute 2x+6 for y in3x−y=6: 3x−y=6 3x−(2x+6)=6 x−6=6(Simplify both sides of the equation) x−6+6=6+6(Add 6 to both sides) x=12 Step 3: Substitute 12for x iny=2x+6: y=2x+6 y=(2)(12)+6 y=30(Simplify both sides of the equation)

OpenStudy (zombrexi):

So now, what do you think the answer is?

OpenStudy (anonymous):

y=60+6???

OpenStudy (anonymous):

im having a a major brainfart im really tired sorry

OpenStudy (zombrexi):

Nope. Y=30 and X=12

OpenStudy (anonymous):

thanks so much

OpenStudy (zombrexi):

Haha its okay. We all have our days.

OpenStudy (zombrexi):

No problem!